OpenStudy (anonymous):

980 is what percent of 1754? Round to the nearest hundredth.

OpenStudy (anonymous):

1754 x = 980 x = 980/1754 = 490/877 490/877 x 100 around 55.87 percent

OpenStudy (anonymous):

ty so much I was hopign I had it right:)